### v1

# OncoReg: Medical Image Registration for Oncological Challenges

Abstract
In modern cancer research, the vast volume of medical data generated is often underutilised due to challenges related to patient privacy. The OncoReg Challenge addresses this issue by enabling researchers to develop and validate image registration methods through a two-phase framework that ensures patient privacy while fostering the development of more generalisable AI models. Phase one involves working with a publicly available dataset, while phase two focuses on training models on a private dataset within secure hospital networks. OncoReg builds upon the foundation established by the Learn2Reg Challenge by incorporating the registration of interventional cone-beam computed tomography (CBCT) with standard planning fan-beam CT (FBCT) images in radiotherapy. Accurate image registration is crucial in oncology, particularly for dynamic treatment adjustments in image-guided radiotherapy, where precise alignment is necessary to minimise radiation exposure to healthy tissues while effectively targeting tumours.
This work details the methodology and data behind the OncoReg Challenge and provides a comprehensive analysis of the competition entries and results. Findings reveal that feature extraction plays a pivotal role in this registration task. A new method emerging from this challenge demonstrated its versatility, while established approaches continue to perform comparably to newer techniques. Both deep learning and classical approaches still play significant roles in image registration, with the combination of methods—particularly in feature extraction—proving most effective.
Keywords: Machine Learning, Image Registration, Cone-beam Computed Tomography, Image Guided Radiotherapy
1 Introduction
In radiology, and particularly in modern cancer research, vast amounts of data are continuously generated. However, much of this data remains underutilized for developing innovative AI-centered solutions due to several challenges. One significant obstacle is the restriction on data publication imposed by patient privacy rights. This is especially problematic for data collected in previous years, where obtaining retroactive patient consent is no longer feasible. To unlock the potential of such data while ensuring compliance with privacy laws, novel approaches are required.
The OncoReg Challenge addresses this issue by enabling researchers to develop and validate methods in a two-phase framework. In the first phase, challenge participants work with a publicly available dataset to design and refine their algorithms. In the second phase, the submitted models are trained in-house on a private, unpublished dataset—ensuring that the data remains securely within the hospital network and never becomes accessible to external participants.
This approach not only safeguards patient privacy but also fosters the development of more generalizable AI models. Since participants must optimize their methods for unseen data rather than a specific dataset, the challenge encourages the creation of algorithms with broader applicability and better real-world transferability. Computational challenges provide an ideal framework for this kind of data utilization, allowing multiple research groups to benefit from the dataset without the need for public distribution. The popularity of challenges, especially in the field of computer vision, has been steadily increasing, as reflected in the growing number of challenges hosted at major conferences such as MICCAI, ISBI, and SPIE Medical Imaging.
In the field of medical image registration, the Learn2Reg Challenge (Hering et al., 2022) holds a prime position. The well-known challenge, which has been held annually since 2019 in conjunction with MICCAI, has significantly advanced algorithmic approaches and benchmarking in medical image registration, fostering collaboration and innovation among researchers worldwide. Learn2Reg offers an extensive selection of 3D image registration tasks, covering a variety of scenarios. The tasks include both mono- and multimodal registration as well as inter- and intrapatient registration. The challenges span multiple imaging modalities such as computed tomography (CT), magnetic resonance imaging (MRI), and ultrasound (US), and encompass different anatomical regions, including the lungs, abdomen, and brain. OncoReg builds upon the foundation established by Learn2Reg, extending it to include an additional modality—cone-beam CT—and shifting the focus to real-world oncological challenges. The first part of OncoReg comprises a task (ThoraxCBCT) in the Learn2Reg Challenge 2023, followed by a standalone virtual challenge and workshop.
This focus on real-world scenarios highlights the importance of image registration, which has numerous medical applications but plays an especially critical role in oncology, particularly in the context of radiotherapy, where precise targeting of cancerous tissues is vital for effective treatment. Radiotherapy involves delivering high doses of radiation to tumour sites while minimising exposure to surrounding healthy tissues. To achieve this level of precision, accurate alignment of images from different time points and imaging modalities is crucial.
In particular, the registration of low-dose interventional cone-beam CTs (CBCTs) with high-resolution diagnostic fan-beam CTs (FBCTs) is a fundamental step in image-guided radiotherapy. FBCTs, acquired before treatment, provide a detailed map of the patient’s anatomy and are used to plan the radiation dose distribution. However, during the course of treatment, the patient’s anatomy can change due to tumour shrinkage, weight loss, or other factors. CBCTs are frequently acquired during treatment sessions to assess these changes in real time. By registering CBCTs with the initial planning CTs, clinicians can adjust the treatment plan dynamically, ensuring that radiation continues to be delivered precisely to the target area, even as the anatomy evolves.
This registration process is technically challenging due to differences in image quality, patient positioning, and anatomical changes over time. Therefore, developing robust algorithms for accurate image registration is essential to improving the efficacy of radiotherapy. Standardising planning processes and continuously enhancing registration techniques, particularly with the support of adaptive algorithms, not only enhances patient safety but also minimises treatment-related side effects. The development and refinement of such algorithms through challenges like OncoReg contribute to optimising treatment options for cancer patients.
This work presents the methodological approach used in conducting the OncoReg challenge and provides a detailed description and analysis of the entries, results, and rankings.
2 Related Work
2.1 Image Registration Challenges
The EMPIRE10 challenge (Murphy et al., 2011) was one of the first widely recognized image registration challenges, specifically targeting the elastic nature of lung tissue deformations. The challenge evaluated registration performance on 30 thoracic CT scan pairs using several metrics: distances between manual landmark pairs, fissure segmentations, and Jacobian determinant values of the deformation field. A unique feature of the original 2010 workshop was the requirement for live computations, placing time constraints on the registration algorithms. EMPIRE10 set a foundational precedent for evaluating pulmonary image registration and established performance metrics that have influenced subsequent challenges.
Several years later, the Continuous Registration Challenge (Marstal et al., 2019) introduced a continuous benchmarking system with automatic evaluation based on eight datasets, focusing on lung CT and brain MRI registration. However, the framework was somewhat limited by the requirement that algorithms had to be compatible with the SuperElastix platform. In 2019, two other major challenges emerged alongside the first iteration of the Learn2Reg Challenge (Hering et al., 2022). The CuRIOUS challenge (Xiao et al., 2019) addressed brain shift correction using intraoperative ultrasound and preoperative MRI, while the ANHIR challenge (Borovec et al., 2020) focused on registering histological images. In the following years, BraTS-Reg (Baheti et al., 2021) provided a benchmark for registering preoperative and follow-up MRI scans in diffuse glioma patients, and the Fet-Reg challenge (Bano et al., 2021) centered on fetal MRI registration. More recently, the ACROBAT challenge (Weitz et al., 2024) aimed to improve the alignment of stained whole-slide images from breast cancer tissue sections.
While many registration challenges have focused on specific tasks or imaging modalities, Learn2Reg stands out for its comprehensive approach. It covers a wide range of tasks, including mono- and multimodal registration, across diverse anatomical regions and imaging modalities such as CT, MRI, and ultrasound. With the addition of OncoReg to Learn2Reg in 2023 in the form of the ThoraxCBCT task, the challenge introduced its first oncological lung registration task. This also marked the inclusion of cone-beam computed tomography (CBCT) alongside conventional diagnostic fan-beam CT (FBCT), expanding the scope of the challenge to address real-world clinical problems in oncology.
2.2 Cone-Beam CT Registration
The registration of cone-beam CTs has various applications in medicine, with image-guided radiotherapy being one of the most prominent (Zhang et al., 2024; Cao et al., 2022; Han et al., 2021). A related application is image-guided surgery, with a more specialized case being video-assisted thoracoscopic surgery (VATS) (Alvarez et al., 2021), a minimally invasive technique for lung nodule resection in early-stage lung cancer. Beyond oncology, CBCT registration is also utilised in dental imaging (Zheng et al., 2025) and robotic cochlear implant surgery (Lu et al., 2022).
Despite the widespread use of CBCT, its registration to conventional CT remains a challenging problem due to lower image quality, artifacts, and intensity variations. While general-purpose registration methods have been evaluated on CBCT-CT tasks (Demir et al., 2024; van Harten et al., 2024), dedicated solutions remain relatively scarce. Historically, semi-automatic optimization-based methods have been applied (Sotiras et al., 2013), and classical optimization approaches continue to be refined (Xie et al., 2022). More recently, deep learning-based methods have emerged, such as an unsupervised dual-attention network (Hu et al., 2022) and a rigid deep-learning-based registration method for head-neck CT-CBCT (Peng et al., 2024). Additionally, (Teng et al., 2021) introduced a supervised CNN for phase-to-phase registration in 4D CBCT, and (Li et al., 2023) proposed generating high-quality synthetic CTs from CBCTs as an alternative approach to facilitate registration.
While progress has been made, CBCT-CT registration remains an open research challenge, with ongoing efforts to develop more robust, artifact-resistant, and anatomically aware methods that can improve clinical workflows and treatment outcomes.
3 Methods
3.1 Challenge Organisation
The OncoReg challenge was split in two phases, the first of which was organised as part of the 2023 version of the Learn2Reg challenge associated with MICCAI 2023 and the second as MICCAI endorsed virtual workshop in February 2024. The Learn2Reg challenge was carried out as a type 1 and type 2 challenge whereas the second phase resulting in the OncoReg workshop was carried out as a type 3 challenge. The OncoReg challenge was organised by Mattias Heinrich, Alessa Hering, Christoph Großbröhmer and Wiebke Heyer.
Computational challenges in the field of machine learning are often categorised based on the accessibility of data to participants. These challenges generally fall into four distinct types. Type 0 challenges offer participants access to all data phases, including supervision data, which minimises the workload for organisers. However, this approach compromises fairness, as participants are responsible for computing their own metrics, leading to potential inconsistencies. Type 1 challenges improve on this by keeping the supervision data for the test phase private, while the rest of the data is publicly accessible. Participants submit their predictions, which the organisers then evaluate and rank. This structure ensures a more consistent evaluation process but still allows participants access to much of the data. Type 2 challenges further increase the complexity by withholding not just the supervision data but also the test phase data itself. Participants must submit an inference algorithm, which the organisers run on the hidden data before evaluating and ranking the results. This approach increases the workload for both organisers and participants. Participants often need to containerise their methods or compile them into specific formats, adding to the complexity. Type 3 challenges represent the most complex scenario, where both the training and test data are hidden. Participants must submit their entire training algorithm, which organisers then use to retrain models on the hidden data before conducting inference, evaluation, and ranking. While this increases the burden on organisers significantly, it offers several advantages. One key benefit is that sensitive data—such as medical data that cannot be publicly disclosed due to patient privacy concerns—can still be utilised for research and algorithm development. Additionally, this approach ensures that all submissions are evaluated under identical conditions, eliminating the use of additional data, annotations, or varied hardware, thereby drastically enhancing fairness.
Furthermore, keeping data private in this manner promotes the generalisability of solutions. Without access to the training data, algorithms cannot be overfitted to the specific task, potentially leading to more robust models that are applicable across a broader range of scenarios.
3.1.1 Learn2Reg 2023 ThoraxCBCT Task
In 2023, the Learn2Reg challenge expanded its scope by introducing the new ThoraxCBCT task and extending the National Lung Screening Trial (NLST) task from the 2022. It also continued previous tasks such as LungCT, OASIS, and AbdomenMRCT, building on the challenge’s comprehensive history (Hering et al., 2022). The two primary tasks of the competition were hosted entirely on grand-challenge.org and followed a structured, two-phase process:
  • Validation Phase: During this phase, participants developed their algorithms using the provided training and validation data. For evaluation, they were required to compute displacement fields for the validation cases and upload them to grand-challenge.org. Participants were allowed multiple submissions per day, with results and rankings updated in real-time on a public leaderboard. This encouraged iterative improvement and competition among the teams.
  • Test Phase: Participants could either request access to the test data, submitting their computed displacement fields as in a traditional Type 1 challenge, or they could opt to submit their containerised algorithms through the grand-challenge.org/algorithms platform, marking a Type 2 challenge. This second option, which promoted algorithm submissions, was highly encouraged by the organisers to support the publication of user-friendly, ready-to-use methods on the platform. An additional evaluation of the inference runtime was conducted for those submitting containerised algorithms, emphasizing the importance of both accuracy and efficiency in real-world applications.
3.1.2 OncoReg Workshop
The second phase of OncoReg expanded on the Learn2Reg challenge, building upon the ThoraxCBCT task and transforming it into a type 3 challenge that utilised a completely hidden, unpublished dataset. The OncoReg workshop placed a special emphasis on the role of registration within the context of radiotherapy.
Participants were encouraged to develop new methods or adapt their Learn2Reg submissions for the type 3 task, using the ThoraxCBCT leaderboard for initial validation. They were then given the opportunity to submit their containerised solutions for a preliminary sanity check. For deep learning-based approaches, final submissions were retrained and evaluated on the hidden dataset. Due to privacy restrictions on the dataset, this phase of the challenge was not conducted on grand-challenge.org.
3.1.3 Lowering Entry Barriers
To keep the challenge accessible to potential participants with little experience in image registration and containerisation diverse approaches for assistance were pursued. The different steps of participation were accompanied in several ways: To help participants in getting started in their algorithm development pre-processed data was provided and baseline methods for comparison were made public on the leaderboards. Code for evaluation including multiple metrics was published and examples for the creation of an algorithm on grand-challenge.org were provided on GitHub and grand-challenge.org. Regarding the type 3 challenge an example submission docker with step-by-step instructions was provided and sanity checks were offered. Additionally, participants had the opportunity to ask questions and discuss in a public Q&A session, via a discussion forum and via direct email to the organisers. All additional resources for Learn2Reg 2023 and OncoReg can be found on our grand-challenge website [^1] https://learn2reg.grand-challenge.org/oncoreg/ and our GitHub [^2] https://github.com/MDL-UzL/L2R/, https://github.com/MDL-UzL/OncoReg/.
3.2 Tasks & Data
3.2.1 ThoraxCBCT
With the release of the ThoraxCBCT dataset the registration problem of image guided radiation therapy between pre-therapeutic fan-beam CT (FBCT) and interventional low-dose cone-beam CT (CBCT) is addressed. The released image data is part of the 4D-Lung dataset (Hugo et al., 2016) from The Cancer Imaging Archive which contains four-dimensional lung images acquired during radiochemotherapy of locally-advanced, non-small cell lung cancer (NSCLC) patients.
For each patient one FBCT prior to therapy in maximum inspiration and two CBCTs, one at the beginning of therapy and one at the end of therapy, both in maximum expiration, are provided. The task is to find a solution for the registration of two pairs of images per patient:
  1. 1.
    The planning FBCT (inspiration) prior to therapy and the low dose CBCT at the beginning of therapy (expiration), which are acquired at similar timepoints, and
  2. 2.
    The planning FBCT (inspiration) prior to therapy and the low dose CBCT at the end of therapy (expiration), where longer periods of time (usually several months) lay in between the scans.
The registration challenge in both image pairs involves aligning two different modalities: Fan-Beam CT and Cone-Beam CT, while accounting for shifts between maximum inspiration and expiration phases of breathing. The second subtask presents an additional challenge with temporal differences, as it compares the planning CT taken at the start of therapy with the follow-up CBCT acquired at the end of therapy.
The dataset released includes training images from 11 patients, comprising 11 FBCT and 22 CBCT scans, and validation images from 3 patients, which consist of 3 FBCT and 6 CBCT scans. Each patient study includes two image pairs, resulting in a total of 22 training pairs and 6 validation pairs. For every image pair, Förstner keypoints are provided, based on displacement fields computed using ConvexAdam (Siebert et al., 2021). Unlike previous Learn2Reg tasks that focused primarily on the lungs, this task extends to include other thoracic structures, particularly organs at risk in radiotherapy. Keypoints were therefore determined not only within the lung but across the entire trunk region.
The test dataset includes images from 5 patients, yielding 5 FBCT and 10 CBCT scans (10 test pairs). The validation and test annotations feature multiple segmentation masks for various structures, including lungs, tumors, organs at risk (heart, spinal cord, esophagus, trachea, and aorta), and skeletal structures such as vertebrae, ribs, clavicles, and scapulae. Additionally, anatomical landmarks like airway bifurcations, the top of the aortic arch, apex cordis, and key points in bone structures were marked. Some annotations for FBCTs were sourced from TCIA, with missing segmentation masks generated using the TotalSegmentator tool (Wasserthal et al., 2023) and propagated to CBCTs through registration. These generated masks were validated by clinicians at the University Hospital of Schleswig-Holstein (UKSH). On average, 13 structures were segmented per image. Landmark annotations were entirely manual, resulting in an average of 16 landmarks per image pair. A binary body-trunk mask, distinguishing the thorax from the background, was provided for each image, also created using the TotalSegmentator.
All images were converted from DICOM to NIfTI format and subsequently resampled and cropped to the thoracic region, resulting in an image size of 390x280x300 with a voxel spacing of 1.0x1.0x1.0 mm. Large translations between image pairs were manually corrected prior to registration.
Training and validation images, trunk masks and keypoints are available via our grand-challenge website [^3] https://learn2reg.grand-challenge.org/oncoreg/ and test images as well as all supervision data is available upon request or via automatic evaluation on grand-challenge.org.
3.2.2 OncoReg
The OncoReg task extends the ThoraxCBCT challenge but is conducted as a type 3 task using a completely hidden dataset. Participants are tasked with registering a high-resolution diagnostic FBCT acquired before therapy with two CBCT scans: one from the start and another from the end of therapy. Unlike the ThoraxCBCT data, OncoReg images are not synchronized with respect to breathing phases, meaning the scans may capture different states of inspiration and expiration. The data, collected from routine clinical care at the University Hospital of Schleswig-Holstein (UKSH), was retrospectively curated from 25 lung cancer patients and carefully annotated with semi-automatic segmentations and manual landmarks. Similar to ThoraxCBCT, trunk masks and Förstner keypoints were computed for each pair. Preprocessing included converting images from DICOM to NIfTI format, cropping, resampling to a size of 256x192x192, with a voxel spacing of 1.5x1.5x1.5 mm, and manual pre-alignment of image pairs. Annotations for OncoReg were more comprehensive than those for ThoraxCBCT, as the entire dataset was annotated, rather than just a subset. Segmentations were created using the TotalSegmentator tool and then manually refined, resulting in an average of 52 annotated structures per image. Manual landmark annotation was also more detailed, with an average of 25 landmarks identified per image pair. While this dataset is not publicly available, researchers may request evaluation of their registration methods on the data via the challenge platform.
3.3 Challenge Design
To ensure a comprehensive evaluation of the registration performance, a range of complementary metrics are employed that are adapted after previous editions of Learn2Reg and are outlined in Table 1. These metrics are chosen to evaluate various aspects of algorithmic performance, including accuracy, robustness, plausibility, and computational efficiency. The final ranking also accounts for the statistical significance of the differences between results, using a ranking system inspired by the Medical Decathlon (Antonelli et al., 2022). Methods are statistically compared using the Wilcoxon signed-rank test, with a significance level set at $p<0.05$. Based on the number of “won” comparisons, each method receives a numerical rank score ranging from 0.1 to 1. The overall task ranking is then determined by calculating the geometric mean of these individual metric scores.
Table 1: Summary of Metrics used throughout the challenge for evaluation of registration performance.
Metric Description Purpose Unit
DSC Dice Similarity Coefficient Overlap between segmentations %
DSC30 30th Percentile of DSC Measures robustness of overlap %
HD95 95th Percentile Hausdorff Dist. Robust surface distance mm
TRE Target Registration Error Landmark misalignment mm
TRE30 30th Percentile of TRE Robustness of landmark error mm
SDlogJ Standard Deviation of log(Jac.) Field smoothness -
RT Runtime Time taken for registration s
3.4 Challenge Entries
The ThoraxCBCT validation leaderboard featured 68 entries from seven teams (including the challenge organisers), with seven valid submissions made during the test phase. Four of these submissions were baseline methods provided by the organisers.
The type 3 OncoReg task received five submissions from four participating teams, in addition to five baseline methods supplied by the organisers. The baseline methods included Niftyreg (Modat et al., 2009), DeedsBCV (Heinrich et al., 2013, 2015), ConvexAdam (Siebert et al., 2022; Heinrich et al., 2014), and Voxelmorph++ (Heinrich and Hansen, 2022). Voxelmorph++, in particular, was made available as an algorithm on grand-challenge.org/algorithms and was containerised as an example on the challenge’s GitHub repository, which is why it is discussed in detail in the following section. Additionally, a keypoint based registration method using the medical open network for AI (MONAI) (Cardoso et al., 2022) is provided. The baseline methods will not be further elaborated upon here, as they are already well-documented in existing literature. The methods submitted by the participating teams are described in detail below.
3.4.1 DINO-Reg
DINO-Reg (Song et al., 2024) is an optimisation-based method for medical image registration. The process begins with the encoding of medical image features using DINOv2 (Oquab et al., 2024), followed by an optimisation step applied to these encoded features to achieve accurate registration. DINOv2 is a 2D vision transformer pre-trained on natural images without labels, using knowledge distillation. The model demonstrates a strong understanding of image semantics and exceptional generalisability, making it suitable for medical image feature encoding without the need for fine-tuning. To adapt the model for 3D medical image feature extraction, encoding is performed on axial view slices, and the most relevant principal components are extracted from all features within the 3D volume. The resulting features facilitate accurate registration via simple gradient descent optimisation.
Specifically, each 2D slice from the image is encoded into a 3D feature map using DINOv2, with the third dimension representing the feature space. Stacking the 3D feature maps from all slices generates a 4D feature map of the entire volume. Once the 4D feature maps for both the moving and reference images are obtained, principal component analysis (PCA) is performed on all feature tokens to (1) reduce feature dimensionality, and (2) align the moving and reference image features into the same feature space. The principal components aim to capture the variance within the original features. For these features to be meaningful in the context of medical image registration, the principal components must describe anatomical differences, such as those between organs and bones. If left unprocessed, most of the variance would arise between foreground and background content. Therefore, thresholding is applied to the original image intensity to isolate foreground patch features.
Due to the patch-encoding step of vision transformers, the resulting feature map has a resolution 14 times lower than the original image. To mitigate this effect, the original image is up-sampled prior to encoding. During implementation, feature interpolation between slices is applied, and PCA is replaced with low-rank PCA to accelerate the method. For the CBCT data, DINO-Reg features and MIND features are found to be complementary, with the former providing semantic information and the latter capturing edge details. Consequently, a second round of optimisation using MIND (Heinrich et al., 2012a) features was tested after performing DINO-Reg, resulting in a method named DINO-RegEnsemble.
3.4.2 Fourier-Net
U-Net-based methods have dominated medical image registration due to their fast inference speed. The deep models take the full-resolution images ($I_{M},I_{F}$) as input and estimate full-resolution displacements ($\Phi$). To reduce the repeated convolution operations in the U-Net architecture and boost computational efficiency, a model-driven method called Fourier-Net is proposed to learn a low-dimensional representation of the displacement.
Fourier-Net operates under the hypothesis that the desired displacement is smooth, allowing it to be represented using only the low-frequency components. However, it has been experimentally demonstrated in (Jia et al., 2023) that learning the complex-valued frequencies of displacement from spatial images is challenging for a convolutional network. Fourier-Net instead learns a small spatial patch ($\mathbb{S}_{\phi}$) that shares the same low frequencies with the desired full displacement. Theoretically, once the small patch is perfectly learned by the network, incorporating the FFT transform, zero-padding, and inverse FFT transform will get the final displacement:
$\Phi=iFFT(Pad(FFT(\mathbb{S}_{\phi})))$ (1)
However, Fourier-Net only learns the low-frequency components of the displacement, therefore deformation fields estimated might lack local details, limiting its applications to non-smooth and complex registrations. In this challenge, the Fourier-Net employed consists of 5 downsampling convolutional blocks in the contracting path and 1 upsampling convolutional block in the expansive path. Specifically, for an input of size $H\times W\times D$, the resolution after each block progresses as follows: $[C,D,H,W]$, $[2C,\frac{D}{2},\frac{H}{2},\frac{W}{2}]$, $[4C,\frac{D}{4},\frac{H}{4},\frac{W}{4}]$, $[8C,\frac{D}{8},\frac{H}{8},\frac{W}{8}]$, $[16C,\frac{D}{16},\frac{H}{16},\frac{W}{16}]$, and $[3,\frac{D}{8},\frac{H}{8},\frac{W}{8}]$. The variable $C$, representing the number of channels in the first convolutional layer, is set to 16 for this challenge. The Fourier-Net code is available at https://github.com/xi-jia/Fourier-Net.
A pre-trained model from the ThoraxCBCT dataset was utilised to initialise the weights before the training phase. Once the model was trained, instance-level optimisation was employed for each test pair to refine the displacement estimation. Normalised cross-correlation (NCC) and bending energy were used as the similarity measure and regularisation term, respectively. Additionally, a Dice loss on the keypoint masks was incorporated as extra supervision during the network updates.
Keypoint Masks: Instead of using individual key points, each key point was empirically rounded into a $3\times 3\times 3$ patch, and the Dice loss was computed. This approach was chosen because key points, often labelled by experts or classical methods, may contain some degree of subjectivity or error. Rounding the key points into a patch provides a tolerance for such uncertainty.
Image Augmentation: Online data augmentation was applied during pre-training on the ThoraxCBCT dataset. Specifically, images were randomly rotated, cropped, and subjected to noise addition, using the default probability settings provided by https://github.com/ZFTurbo/volumentations. The training process on the private dataset was similar to that used for the ThoraxCBCT dataset, where image similarity, keypoint mask Dice, and bending energy were optimised. Online data augmentation was also employed during this stage. The instance-level optimisation involved fine-tuning the trained weights on each test pair. For this challenge, the number of iterations was fixed at 700, with a learning rate of 0.0001 for each test pair.
3.4.3 SynDeeds
The OncoReg task focuses on the challenging yet clinically crucial objective of aligning pre-therapeutic, high-quality fan-beam CT (FBCT) with interventional low-dose cone-beam CT (CBCT) for image-guided radiation therapy. The low contrast resolution and frequent presence of metal artifacts in CBCT images create significant difficulties in achieving accurate registration.
To address the registration challenges posed by low-quality CBCT images, SynDeeds, an enhanced version of Deeds (Heinrich et al., 2012b), is introduced. This variation incorporates a learning-based image synthesis module (Chen et al., 2019), aiming to enhance the accuracy of registration significantly.
Figure 1: SynDeeds: Fixed and moving images are registered using Deeds after the fixed image went through the images restoration module.
The image synthesis approach is framed as an additive restoration problem, defining the clean image as the sum of a degraded image and a variable “noise” component. The loss from network learning is computed as the mean squared error (MSE) between the ground truth residual and the output residual. Therefore, the image synthesis network is trained to learn the residual, instead of reconstructing the clean image directly, as shown below:
Figure 2: Overview of the image restoration module.
3.4.4 Voxelmorph++
Voxelmorph++ (Heinrich and Hansen, 2022) extends the capabilities of the well-known medical image registration method Voxelmorph (Balakrishnan et al., 2019) by incorporating a novel heatmap prediction strategy with keypoint-based self-supervision as well as multi-channel instance optimisation. Unlike Voxelmorph, which may be susceptible to ambiguous optimisation landscapes due to the spatial transformer loss, the novel heatmap approach aims to capture multiple scales of deformations within a single feed-forward network.
Voxelmorph++ utilises a basic single-scale U-Net (Ronneberger et al., 2015) architecture as its baseline backbone, processing image pairs to extract a shared feature representation. This representation serves as the foundation for predicting displacement fields that aim to align fixed and moving images within a designated region of interest.
A distinctive feature of Voxelmorph++ is its discretised heatmap prediction head, which translates the shared feature vectors into a probabilistic tensor representing the likelihood of displacements at specific keypoints. Leveraging pre-generated automatic correspondences based on Förstner keypoints for self-supervision, the prediction head robustly reduces large deformations and enables discretised integral regression of displacements to directly and explicitly match the keypoint supervision. The heatmap is predicted using a transpose 3D convolution followed by several 3D convolution blocks. The output heatmap tensor is then used to define probabilistic displacements, facilitating a more precise alignment by accounting for local variations and potential discrepancies in displacement estimation.
The addition of Förstner keypoints as a part of the self-supervision strategy enhances the method’s ability to capture distinctive features within the images, improving the accuracy of the registration process. These keypoints serve as reference points for guiding the model’s predictions, resulting in more precise alignment between the fixed and moving images. In case no keypoint correspondences are available heatmaps can still be utilised in an unsupervised matter using only the non-local MIND metric [2] and Förstner keypoints in the fixed scan to focus on distinct locations within the region of interest.
For the OncoReg task Förstner keypoints were generated for the moving image and ConvexAdam (Siebert et al., 2021) was used to establish automatic correspondences by registering the moving image to the fixed image and applying the transformation to the keypoints. Keypoints were generated for a masked image region covering the patients trunk. Using this method does not result in ground-truth comparable correspondences but it does provide a weak supervision, which proved to work sufficient for the task. Additionally, the challenge implementation of the method makes use of the provided trunk masks to limit the region of interest and applies affine augmentation to both, the masked images and the keypoints.
Voxelmorph++ employs a non-local MIND loss function, which enhances the model’s ability to handle the inherent ambiguities and complexities of medical image registration. This loss function computes by comparing warped MIND descriptors of the moving scan, weighted by the predicted displacement probabilities, thus avoiding the pitfalls associated with direct spatial transformations. The predicted sparse displacements are extrapolated to a dense displacement field using thin-plate-splines.
The framework also integrates instance optimisation, which replaces multiple learned fine-tuning steps and refines the initial predictions through iterative corrections using a robust regularisation strategy comprising a combined B-spline and diffusion regularisation coupled with a MIND metric loss. This optimisation step enhances the accuracy of the registration and ensures robustness against local minima. An ablation study proved this step to be imperative for high performance on the OncoReg task.
In summary, Voxelmorph++ achieves state-of-the-art performance in large-deformation registration by combining a single-scale U-Net backbone, heatmap prediction, non-local MIND loss, and instance optimisation. This approach provides a robust and efficient framework for medical image registration tasks, with applications in both intra-subject and inter-subject studies.
3.4.5 SSKJLBW: Unsupervised Single-shot Registration
To address the challenges of severe noise interference, large-scale deformations caused by respiratory motion, and limited training data (only 10 cases) in thoracic cone-beam CT (ThoraxCBCT) image registration for clinical radiotherapy, this paper proposes a single-shot unsupervised registration framework. Inheriting the optimization paradigm of GroupRegNet (Zhang et al., 2021), the method achieves robust registration without annotated data through multi-scale feature fusion and dynamic regularization constraints. The core workflow begins with noise-adaptive preprocessing of input images, followed by iterative optimization of deformation fields via a cascaded estimation module, and finally outputs smoothed deformation fields constrained by anatomical masks.
The framework incorporates physics-guided preprocessing to decouple noise from anatomical deformations. Input HU values are clamped to [0,2048] to suppress metal artifacts and extreme noise, while edge regions are zero-padded to mitigate respiratory motion artifacts. A lightweight 3D UNet processes standardized 128×128×64 image pairs, leveraging hierarchical features for deformation estimation—shallow layers enforce smooth large-scale deformations, while deeper layers refine local details.
Central to the approach is a dynamic template evolution strategy. Initial reference templates are generated by applying random affine transformations to moving images. During iterative alignment, templates are updated by blending historical and newly aligned results, with blending weights gradually shifting from stability-focused (75% historical) to sensitivity-focused (80% new) across iterations. This balances global consistency and local precision.
Training leverages anatomy-aware augmentation within torso masks derived from TotalSegmentator (Wasserthal et al., 2023), combining spatial transforms (rotation, scaling, translation) with synthetic noise injection. The loss function integrates noise-robust local cross-correlation for structural alignment, spatial gradient penalties for smooth deformations (Liu et al., 2020), and Jacobian regularization to prevent topological folding.
For inference, a coarse-to-fine hierarchical strategy accelerates computation. Low-resolution (1/8) coarse alignment initializes deformation fields, which are progressively refined at 1/2 and full resolutions. Postprocessing with median filtering removes outlier deformations, achieving clinically feasible 3-minute runtime per case. Experimental validation demonstrates 0.84 Dice for lung alignment under extreme noise (SNR¡5dB), outperforming traditional methods in accuracy and efficiency. The method’s noise-deformation decoupling and adaptive template updating provide a practical solution for limited-data scenarios in thoracic image registration.
4 Results
4.1 ThoraxCBCT
The quantitative results for the ThoraxCBCT task are summarized in Table 2. While ConvexAdam achieved the highest overall ranking across all evaluation metrics, the official winner of the task was SynDeeds. Since ConvexAdam was provided by the organizers as a baseline method, it was not eligible for ranking. However, the performance of ConvexAdam and SynDeeds was very close across most metrics.
When examining individual performance measures, SynDeeds achieved the highest mean Dice Score, while Fourier-Net obtained the lowest Target Registration Error (TRE). The best Hausdorff Distance 95th percentile (HD95) was observed for SSKJLBW, although this method exhibited a less smooth displacement field.
A more detailed breakdown of Dice coefficients achieved by all methods across different anatomical labels is shown in Fig. 3, while the distribution of Target Registration Errors (TREs) is presented in Fig. 5.
4.2 OncoReg
For the initial challenge evaluation, we used a standard train-validation-test split. However, since many submitted methods did not rely on training data, we later expanded our evaluation to be more robust and meaningful. After completing additional annotations for all images, we conducted a three-fold cross-validation for all trainable models to ensure a fair and comprehensive comparison.
All trainable models were trained and evaluated on an NVIDIA RTX 4000 SFF Ada Generation GPU with 20 GB of memory. Importantly, runtime was not considered in the ranking.
The quantitative results are summarized in Table LABEL:tab:oncoreg and visualized in Fig.6 and Fig.5.
The challenge was won by the DINO-Reg Ensemble, which outperformed all other methods by achieving the lowest Target Registration Error (TRE) and the highest Dice score. It was closely followed by Voxelmorph++ and ConvexAdam, while DeedsBCV and NiftyReg ranked lowest.
To ensure transparency in our evaluation process, we provide the complete code for evaluation and ranking on GitHub [^4] https://github.com/MDL-UzL/L2R/.
Table 2: Quantitative Results of the ThoraxCBCT task.
[…]
Initial 11.91 16.26 29.04 48.13
ConvexAdam 4.93 4.98 68.40 38.00 0.10 0.710
SynDeeds 6.01 5.67 70.16 22.97 0.15 0.705
Niftyreg 4.90 4.51 62.11 43.41 0.05 0.577
DeedsBCV 8.76 7.38 69.00 22.38 0.16 0.537
Fourier-Net 4.03 3.85 61.91 47.80 0.07 0.535
SSKJLBW 5.80 7.44 69.45 17.60 0.26 0.515
Voxelmorph++ 13.06 16.75 46.28 33.34 0.12 0.357
Figure 3: ThoraxCBCT: Dice Coefficients for each anatomical label and registration method, averaged across all test images.
Figure 4: ThoraxCBCT: Target Registration Error averaged over all landmarks and test images for each registration method.
Figure 5: OncoReg: Target Registration Error averaged over all landmarks and test images for each registration method.
Table 3: Quantitative Results of the OncoReg task.
[…]
Initial 6.41 11.55 34.38 77.56
DINO-RegEns. 3.94 7.62 63.89 76.59 0.062 $>300$ 0.71
Voxelmorph++ 4.66 8.52 62.36 73.96 0.081 $<60$ 0.66
ConvexAdam 4.16 7.94 60.13 76.54 0.066 $(<5)$ 0.61
SynDeeds 4.51 7.81 59.93 74.85 0.100 x 0.56
DINO-Reg 4.53 8.50 57.64 77.03 0.040 $>300$ 0.54
Fourier-Net 6.01 10.46 60.67 71.91 0.103 $>300$ 0.50
MONAI SegResNet 4.78 8.96 53.63 77.37 0.048 x 0.49
DeedsBCV 7.99 11.78 58.33 71.16 0.154 $<15$ 0.40
Niftyreg 7.98 12.43 37.30 75.15 0.061 $<60$ 0.261
Figure 6: OncoReg: Dice Coefficients for ten anatomical labels and each registration method, averaged across all test images.
5 Discussion
5.1 Evaluation Metrics and Annotation Challenges
Evaluating image registration remains a complex task, as standard metrics often fail to capture all aspects of registration quality and are highly dependent on annotation accuracy. In this challenge, the primary evaluation metric was the Target Registration Error (TRE), which relies on manually placed anatomical landmarks. However, landmark positioning is inherently uncertain in certain structures; while the carina served as a reliable reference point, the aortic arch and vertebral midpoints were more challenging to pinpoint precisely. To minimize inter-observer variability, all landmarks within a given image pair were annotated by the same medical expert, ensuring internal consistency.
Significant effort was dedicated to refining these annotations, including extending the publicly available 4D-Lung dataset and comprehensively annotating our private dataset. This investment considerably enhanced the dataset’s usability. However, annotation challenges were particularly pronounced in cone-beam CT (CBCT) images, where lower resolution, poor contrast, and strong artifacts complicated precise landmark placement. In these cases, tools like TotalSegmentator proved invaluable, significantly reducing annotation time while maintaining quality in a semi-automated setting.
5.2 Challenge Design and Implementation
Participants had the option to submit results via an automatic evaluation platform on Grand Challenge for the ThoraxCBCT task, but this feature was rarely utilized. Instead, most teams preferred local evaluation. In contrast, the provided Docker example was widely adopted, particularly for handling data loading and saving, which streamlined debugging in the Type 3 challenge setup. Despite these advantages, some recurring issues were identified that should be addressed in future Type 3 challenges.
Organizing such a challenge demands substantial time and resources, but efficient communication and accessibility between organizers and participants played a crucial role in ensuring a smooth workflow. Future iterations should build on this experience to optimize logistical aspects and further reduce participation barriers.
5.3 Key Findings and Methodological Insights
A key takeaway from the challenge is that the best-performing registration methods leveraged strong feature extraction strategies. The top-performing methods—ConvexAdam and DINO-Reg—demonstrate two distinct but effective approaches:
  • ConvexAdam combines handcrafted and semantic features within a two-stage optimization process. The extracted semantic features play a crucial role in achieving robust registration across different modalities.
  • DINO-Reg employs self-supervised learning with Vision Transformers (ViTs), using deep feature representations from a DINO-pretrained ViT to enhance registration performance across modalities.
Both methods prioritize feature extraction but differ in execution—ConvexAdam explicitly designs features, whereas DINO-Reg learns them in a self-supervised manner.
Beyond these top approaches, an intriguing finding was the potential for further improving ConvexAdam’s performance by integrating it with VoxelMorph++. Specifically, using ConvexAdam-derived Förstner keypoints as input for VoxelMorph++ led to a refined registration process. This effect was particularly evident in OncoReg, where preprocessed keypoints significantly improved VoxelMorph++’s accuracy. However, in ThoraxCBCT, VoxelMorph++ performed notably worse than other methods. Further analysis revealed that its performance was hampered by suboptimal keypoint registration, which proved particularly detrimental in the weakly supervised setting. A post-challenge re-evaluation using keypoints registered by ConvexAdam showed substantial improvement—although this did not affect the official rankings.
Another notable result was the enhancement of DeedsBCV by Li et al., leading to the superior performance of SynDeeds in both registration tasks. This demonstrates that traditional methods, when effectively adapted, can still compete with or even outperform newer deep-learning-based approaches.
5.4 Conclusion and Future Directions
The OncoReg challenge provided valuable insights into the current state of medical image registration, particularly for oncological applications involving CBCT-to-FBCT registration. Our findings highlight three key aspects:
  • Feature extraction is critical: Whether through explicitly designed features (ConvexAdam) or learned representations (DINO-Reg), robust feature extraction remains a cornerstone of successful registration.
  • Hybrid approaches show promise: Combining classical registration techniques with deep learning can yield improvements, as seen with ConvexAdam and VoxelMorph++.
  • High-quality annotations remain a challenge: While extensive annotation efforts significantly improved dataset usability, CBCT annotations remain difficult due to image quality limitations.
Future iterations of the challenge should focus on refining evaluation protocols, further improving annotation pipelines, and exploring hybrid registration methods that integrate learning-based and conventional techniques. Additionally, optimizing the challenge’s logistical framework, particularly Type 3 implementations and automated evaluation workflows, could enhance participant engagement and improve the overall assessment process.
Acknowledgments
This work was partly funded by the German Federal Ministry of Education and Research (BMBF) (grant number 01KD2212A).
Ethical Standards
The work follows appropriate ethical standards in conducting research and writing the manuscript, following all applicable laws and regulations regarding treatment of animals or human subjects.
Conflicts of Interest
We declare we don’t have conflicts of interest.
Data availability
The data for the ThoraxCBCT task is available on learn2reg.grand-challenge.org/oncoreg/. As described earlier the data of the OncoReg task can not be made publicly available, nevertheless, we are open for requests about evaluation on this data that will be computed in-house by us.
A Additional Result Overviews
Figure 7: OncoReg: Dice Coefficients for individual vertebrae-labels and each registration method, averaged across all test images.
Figure 8: OncoReg: Dice Coefficients for individual rib-labels and each registration method, averaged across all test images.